OpenStudy (elonasushchik):

The equation of line AB is y = 2x + 4. Write an equation of a line parallel to line AB in slope-intercept form that contains point (3, -2).

Nnesha (nnesha):

what is the slope of this line y = 2x + 4.??

OpenStudy (amorfide):

if a line is parallel, then it has the same slope as the line it is parallel to then you use the equation \[y-y _{1}=m(x-x _{1})\] where \[(3,-2)=(x _{1},y _{1})\]

OpenStudy (elonasushchik):

a.) y = 2 x + 4

OpenStudy (elonasushchik):

b.) y= -1/2 x – ½

OpenStudy (elonasushchik):

c.) y= -1/2 x – 7/2

OpenStudy (elonasushchik):

d.)y= 2x - 8

OpenStudy (elonasushchik):

Those are the choices soo..^^^^^^

Nnesha (nnesha):

y = 2x +4 y = mx + b slope intercept form where m is slope which is always with x and b is y-intercept so what is m in tht equation ?

OpenStudy (elonasushchik):

I think its the second choice...

Nnesha (nnesha):

it can't be the 2nd one parallel lines should have same equation the equation should have same slope like this one y = 2x+ 4 <=== what is the slope ofthis equation

OpenStudy (elonasushchik):

oops.. srry..it's the first one

Nnesha (nnesha):

how ??

OpenStudy (elonasushchik):

im sure its the first..

Nnesha (nnesha):

okay i'm asking how did you get first one ?? answer my question please "y = 2x+ 4 <=== what is the slope ofthis equation" ??

OpenStudy (elonasushchik):

okay soo the slope is Slope 4.000/2.000 = 2.000 ans like u said parallel lines should have the same equations soo i know its not the second one for sure and the third or fourth cuz they are negative soo..that's how i got that answer my little @Nnesha !!

Nnesha (nnesha):

fourth one is not negative how did you get 4000/2000

Nnesha (nnesha):

 m represent slope which is always with x